esions -3 Pely on the following information: Firm B is a monopolist that faces market demand Q 200 -2P. Firm Bs tota 3. What is Firm Bs profit maximizing output level (2)? l cost is given by TC() 20 200 + 200. (Hint: Inverse demand is given by P 100-,so total revenue is TR marginal revenue is MR = 100-Q) 1000-9,sa 4. What is Firm Bs profit maximizing price (P)? 5. How much profit is Firm B earning given this price and output level (from problems 3 and 4)?
Answer #1

question 3 ) Q = 200 - 2P

P = 100 - Q/2 -------1)

TR = P*Q = 100Q - Q2/2

MR = Dtr/dq = 100 - Q

TC = 2Q2 + 20Q + 200

MC = Dtc/dq = 4Q + 20

monopolist equates MR = MC

100 - Q = 4Q + 20

80 = 5Q

Q = 16

QUES 4) PROFIT MAXIMIZING PRICE

PUTTING VLAUE OF Q IN EQUATION 1

P = 100 - 16/2

= 100 - 8

= $92

QUES5 ) PROFIT = TR - TC

16 * 92 - ( 2(16* 16 ) + 20 *16 + 200 )

1472 - ( 512 + 320 + 200 )

1472 - 1032

= $440